You are considering a 15-year, $1,000 par value bond. Its coupon rate is 9%, and interest is paid semiannually.

If you require an "effective" annual interest rate (not a nominal rate) of 8.05%, how much should you be willing to pay for the bond? Do not round intermediate steps. Round your answer to the nearest cent.
Answer #1

Value of Bond = 1 1 (1+r) Coupon * + MaturityValue (1 + r) 7

Where r is the discounting rate of a compounding period i.e. 7.89%/ 2 = 3.945%

EAR = (1+Nominal Rate / 2 )^2 -1

0.0805 = (1+r/2)^2 -1

1.0805^1/2 = (1+r/2)

1+r/2 = 1.03947101931

r = 7.89%

And n is the no of Compounding periods 15 years * 2 = 30

Coupon 9% /2 = 4.5%

= \small 45 *\frac{1-\frac{1}{(1+0.03945)^{30}}}{0.03945} + \frac{1000}{(1+0.03945)^{30}}

= 783.36435071 + 313.25058626

= $1096.61

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.